5. The body with mass m2 = 10 kg slides on the horizontal table with coefficient of friction = 0.2. The mass of the hanging body is m1 = 2.5 kg. The two bodies are connected by a light and elongated thread, which passes through the small scroll fixed to the edge of the table. Will the system slide if m1 = 1.5 kg?
1
Expert's answer
2020-12-17T07:26:26-0500
"(m_1+m_2)a=(m_1-\mu m_2)g\\(m_1+m_2)a=(1.5-0.2(10))g=-1.5g<0"

Answer: no.